OpenStudy (anonymous):

solve x --- x^2=8x-35 pleaseee

OpenStudy (igreen):

Do you mean \(x^2 + 8x - 35\)?

OpenStudy (igreen):

Because if it's \(x^2 = 8x - 35\) 'x' is undefined.

OpenStudy (anonymous):

yeaa i

OpenStudy (anonymous):

imaginary answer ...

OpenStudy (igreen):

You can graph it here: https://www.desmos.com/calculator Find the x-intercept. I got 3.141

OpenStudy (anonymous):

how can u simplified this \[\frac{ -8\pm \sqrt{204} }{ -2 }\]

OpenStudy (anonymous):

i use the quadratic formula

OpenStudy (anonymous):

@phi @ganeshie8 any ideas ???

OpenStudy (igreen):

\(\Large\frac{ -8\pm \sqrt{204} }{ -2 }\) \(\Large\frac{ -8\pm 14.28 }{ -2 }\) THIS BREAKS INTO TWO PARTS! \(\Large\frac{ -8+ 14.28 }{ -2 }\) \(\Large\frac{6.28}{ -2 }\) \(-3.14\) --------------------- \(\Large\frac{ -8- 14.28 }{ -2 }\) \(\Large\frac{ -22.28 }{ -2 }\) \(11.14\)

OpenStudy (dawnr):

4*51 xD

OpenStudy (anonymous):

mhhh ok thank you

OpenStudy (igreen):

Medal? c: @meln_n